# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v1

@@ III Results
In this section we describe inferences about the EoS, both to calibrate our methods against previous work and to identify challenges.
@@ III.1 Inferences using galactic pulsars
Figure 8: EoS inference from galactic pulsars and symmetry energy, $\Gamma$-spectral family: Marginal posterior distribution for the combined inference of observations PSRs J0030, J0348, J0740, J1614, and applying the PREX symmetry energy constraint. Top-left panel: Distribution of hyperparameters. Color scale shows the marginal likelihood, with red indicating the largest values and the solid contour shows the 90% credible interval, while one-dimensional histograms show a one-dimensional marginal distribution. Top-right panel: 90% credible interval of pressure versus density, evaluated at each density. Bottom panel: 90% credible interval for radius evaluated at each NS mass, rendered as mass versus radius. 98% interval of the employed priors (initial grid) are shown in yellow.
@@
Figure 9: Impact of parameterization on EoS inference: Like Figure 8, but adopting the $\Upsilon$-spectral EoS, and constraining for only the PSRs J0030, J0348, J0740, J1614 and PREX symmetry energy. We also provide a benchmark analysis with the same PSRs, and symmetry energy, but includes GW constraint and uses $\Gamma$-spectral parameterization for these constraints [45].
@@
We first reanalyze previously-studied events with our approach, initially using the original spectral parameterizations and priors for the assumed nuclear equation of state. To illustrate how our analysis proceeds using comparable constraints as in [45] (i.e., galactic neutron star observations and the nuclear symmetry energy), the colored scatterplot points in top left panel of Figure 8 show the marginal likelihoods versus four model hyperparameters (here, spectral parameters). In this example, we are employing all the galactic pulsar mass and radius measurements mentioned above (i.e., mass-radius information for PSR-J0030, and mass measurements for J0740, J0348, J1614), as well as the lower bounds on maximum neutron star mass from observed massive pulsars, as indicated in Figure 2. As noted above, these likelihood evaluations have been identified by an iterative exploration of our model space, in the process recovering the posterior distribution, whose 90% credible intervals and one-dimensional marginal distributions are also shown in the same panel of Figure 8. As expected given how broad and weak our present observational constraints on neutron star radius as illustrated in Figure 2, we find our marginal likelihoods are relatively uniform over a wide range of the parameter space, insofar as EoS compatible with those parameters exist. (Most of our nominally uniform prior volume encompasses incompatible EoS.) Next, the solid curves in the center panel of Figure 8 express the 90% credible interval of our inferred equation of state (in $p(\rho)$). Once again, we see that our posterior 90% credible pointwise interval only differs modestly from the full range allowed by our prior assumptions, being somewhat narrower at all densities than the ranges shown in Figure 1. Overall, because our prior assumptions are very well adapted to contemporary theoretical preferences, the new NICER observations add relatively little new additional information over and beyond our original preferences, with one important exception highlighted below.
@@
To investigate the impact of the EoS parameterization (and prior), we also inferred the EoS with the $\Upsilon$-spectral family. Figure 9 shows a comparable analysis derived using the $\Upsilon$- spectral parameterization (e.g. Ref. 3), illustrating the inferred relationships between pressure and density; mass and radius; and sound speed versus density. In this figure we construct $p(\rho)$ posterior distributions using one-dimensional credible intervals at each density, while $M(R)$ is constructed similarly versus $M$. In sharp contrast to our fiducial analysis with the $\Gamma$-spectral parameterization, the $\Upsilon$-spectral parameterization converges to extreme EoS at low and high density. Close to our matching density with the fixed crust EoS, this $\Upsilon$-spectral posterior converges to sharp density transitions, consistent with $c_{s}\simeq 0$. At the densities $\gtrsim 10^{14.5}$ g cm-3, however, $c_{s}\to\infty$. In other words, this alternative parameterization (with broader priors) best fits galactic pulsar observations with strong low-density phase transitions and unexpected high-density behavior, so the $M(R)$ curve aligns more closely with astrophysical results, notably the NICER observation. While the EoS $p(\rho)$ and $c_{s}(\rho)$ have unexpected behavior, this model family of course produces mass-radius behavior consistent with astrophysical constraints, which are most informative in radius over a narrow mass range.
@@ III.2 Adding information from GW sources
Figure 10: GW constraints for different astrophysical priors Scatterplot of $\ln{\cal E}$ versus $R_{1.4}$ evaluated for the same three astrophysical priors for NS binaries illustrated in Figure 5. As in that figure, this representation uses evaluations drawn from our default $\Gamma$-spectral EoS discussed above. Top panel: Likelihood for the fiducial spin and mass ratio priors. The GW signal only places an upper bound on the NS radius, not a lower bound. Center panel: As above, but for the high-mass ratio prior. This prior more strongly and more frequently penalizes higher NS radii, though some configurations with large radii are still highly likely. Bottom panel: As above, but for the comparable-$q$ and positive-$\chi_{i,z}$ prior. Unlike the other astrophysical priors, this marginal likelihood strongly disfavors very small NS radius.
@@
Gravitational wave measurements, currently favoring smaller radii, are by our prior’s construction more informative than NICER radius constraints. As noted above, our EoS family and prior enforce a reasonably high radius for low-mass NS, leading through regularity of the mass-radius curve and a need for large maximum mass to EoS models that favor comparable or larger radii for typical NS. For context, comparing Figures 2 and 1, our prior favors NS radii largely consistent with the radius of J0030, including a wide range between $9{\rm km}$ and $16{\rm km}$ for a fiducial $1.4M_{\odot}$ NS. By contrast, as surveyed in Figure 5, GW observations of GW170817 support typically smaller radii for a similar fiducial NS, with little support above $13{\rm km}$ in a direct interpretation of our fiducial GW analyses. As a result and as seen in previous studies, a joint analysis incorporating both galactic pulsars and GW observations finds fiducial pulsar radii comparable to the intersection of the two limits presented above. Figure 11 shows our fiducial calculation incorporating galactic pulsar measurements, GW170817 (with our fiducial prior), and the symmetry energy, following [45]. Using similar parameterizations and observational constraints as a previous study, we find comparable inferences for the nuclear equation of state [45]. Some small differences between our inferences and prior work are expected for technical reasons; for example, we present 90% quantiles in radius at each fixed NS mass, while previous and other works adopt a NS central density prior distribution; see the appendix for further discussion.
@@
That said, our GW analyses above also demonstrate that hidden astrophysical priors have a substantial impact on the inferred NS radius. To characterize this effect, Figure 10 shows a one-dimensional scatterplot of $\ln{\cal E}$ versus $R_{1.4}$ using the three different GW analyses presented in Figure 5. These three analyses draw on identical RIFT marginal likelihood data, differing only in the assumed binary NS population model for mass ratio and spin. This figure affirms (directly within the context of EoS inference) that hidden astrophysical priors embedded in the GW-only analysis will propagate directly into different conclusions about the nuclear EoS.
@@
The pulsar-only likelihoods represented in Figures 2 and 6 and the (multiple) GW-only likelihoods characterized by Figure 10 show that the pulsar and GW measurements may or may not agree, with the amount of tension depending strongly on hidden astrophysical priors. In one scenario, the pulsar/GW results may be in modest tension, consistent with previously presented work. In another scenario, the two may be in good agreement, favoring radii of $13~{}{\rm km}$, if we assume GW170817 is similar to galactic pulsars. Finally, the two approaches could have strong tension, with HESS/GW favoring small radii and NICER favoring large radii, if we assume the kilonova associated with GW170817 favors asymmetric NS binaries. As a concrete example, Figure 12 shows an analysis comparable to Figure 11, excepting only the mass ratio prior ($q<0.6$) applied for GW170817. Despite adopting the same raw GW observation and inference strategy, the same galactic pulsar measurements, and the same symmetry energy, in this scenario we would infer a notably smaller radius for most pulsars, compared to previous work.
@@
Figure 12: Joint EoS constraint from GW, galactic pulsars, and symmetry energy, derived using $\Gamma$-spectral EoS parameterization, using three different prior assumptions about GW170817. The blue contours show results derived using the conventional 170817 prior (uniform masses); the orange requires $\chi_{i,z}>0$; and the green requires $q<0.6$. The analyses shown in blue have also been presented in Fig. 11 Top-left panel: 90% percentile pointwise confidence intervals for $p(\rho)$. Top-right panel: 90% percentile pointwise confidence interval for radius at each mass. Bottom panel: $\Gamma$-spectral parameter posterior distribution. In the bottom panel, we show marginal likelihood evaluations accumulated for the fiducial analysis (corresponding to the blue contours).

### v2

@@ III Results
In this section we describe inferences about the EoS, both to compare our methods against previous work and to identify challenges, along with demonstrating the impact of NS population distribution on source properties.
@@ III.1 Inferences using galactic pulsars
Figure 5: EoS inference from galactic pulsars and symmetry energy, $\Gamma$-spectral family: Marginal posterior distribution for the combined inference of observations PSRs J0030, J0348, J0740, J1614, and applying the PREX symmetry energy constraint. Top-left panel: Distribution of $\Gamma$-spectral hyperparameters. Color scale shows the marginal likelihood, with red indicating the largest values and the solid contour shows the 90% credible interval, while one-dimensional histograms show a one-dimensional marginal distribution. Top-right panel: 90% credible interval of pressure versus density, evaluated at each density. Bottom panel: 90% credible interval for radius evaluated at each NS mass, rendered as mass versus radius. In the figures in the top-right and bottom the 98% interval of the region covered by the initial samples supplied to HyperPipe is also shown.
@@
Figure 6: Impact of parameterization on EoS inference: Like Figure 5, but adopting the $\Upsilon$-spectral EoS, and constraining for only the PSRs J0030, J0348, J0740, J1614 and PREX symmetry energy. We also provide a benchmark analysis with the same PSRs, and symmetry energy, but includes GW constraint and uses $\Gamma$-spectral parameterization for these constraints for reference [45].
@@
We first reanalyze previously-studied events with our approach, initially using the original spectral parameterizations and priors for the assumed nuclear equation of state. To illustrate how our analysis proceeds using comparable constraints as in [45] (i.e., galactic neutron star observations and the nuclear symmetry energy), the colored scatterplot points in top left panel of Figure 5 show the marginal likelihoods versus the four model hyperparameters ($\Gamma$-spectral parameters in this case). In this example, we are employing all the galactic pulsar mass and radius measurements mentioned in Section II.3 (i.e., mass-radius information for PSR-J0030, and mass measurements for J0740, J0348, J1614), as well as the lower bounds on maximum neutron star mass from observed massive pulsars, as indicated in Figure 2. As described in Section II.7, these likelihood evaluations have been identified by HyperPipe by conducting an iterative exploration of the model space, in the process recovering the posterior distribution, whose 90% credible intervals and one-dimensional marginal distributions are also shown in the same panel of Figure 5. As expected given how broad and weak our present observational constraints on neutron star radius as illustrated in Figure 2, we find our marginal likelihoods are relatively uniform over a wide range of the parameter space, insofar as EoS compatible with those parameters exist. (Most of our nominally uniform prior volume encompasses incompatible EoS.) Next, the curves in the top-right (bottom) panels express the 90% credible interval of our inferred equation of state in $p(\rho)$ ($M(R)$). Once again, we see that our posterior 90% credible pointwise interval only differs modestly from the full range suggested by our initial assumptions, being narrower at all Mass-Radius estimates. The initial sample space is only indicative of the EoSs taken into consideration at first, as HyperPipe selects new points to explore, the posterior is free to go beyond this initial sample limits to attain the best fit. Overall, because our prior assumptions are very well adapted to contemporary theoretical preferences, the new NICER observations add relatively little new additional information over and beyond our original preferences, with one important exception highlighted below.
@@
To investigate the impact of the EoS parameterization (and prior), we also inferred the EoS with the $\Upsilon$-spectral family. Figure 6 shows a comparable analysis derived using the $\Upsilon$- spectral parameterization (e.g. Ref. 3), illustrating the inferred relationships between pressure and density, and mass and radius. In this figure we construct $p(\rho)$ posterior distributions using one-dimensional credible intervals at each density, while $M(R)$ is constructed similarly versus $M$. In sharp contrast to our fiducial analysis with the $\Gamma$-spectral parameterization, the $\Upsilon$-spectral parameterization converges to extreme EoS at low and high density. Close to our matching density with the fixed crust EoS, this $\Upsilon$-spectral posterior converges to sharp density transitions, consistent with $c_{s}\simeq 0$. At the densities $\gtrsim 10^{14.5}$ g cm-3, however, $c_{s}\to\infty$. In other words, this alternative parameterization (with broader priors) best fits galactic pulsar observations with strong low-density phase transitions and unexpected high-density behavior, so the $M(R)$ curve aligns more closely with astrophysical results, notably the NICER observation. While the EoS $p(\rho)$ and $c_{s}(\rho)$ have unexpected behavior, this model family of course produces mass-radius behavior consistent with astrophysical constraints, which are most informative in radius over a narrow mass range.
@@ III.2 Adding information from GW sources
In order to jointly infer the nuclear EOS simultaneously using PSR and GW observations, we need the gravitational wave likelihood $L(X)$ in Eq. (9), to compute the appropriate marginal evidence ${\cal E}$ for each EOS. In this section, we first describe our reanalysis of GW170817, with which we construct the necessary marginal likelihood. We then demonstrate, using this phenomenological approach alone and without recourse to detailed EOS modeling, that the posterior for $\tilde{\Lambda}$ depends strongly on the assumed BNS population prior adopted, specifically using three concrete examples motivated in Section II.4. We discuss the correlations and features of the GW posterior that cause these prior changes to have the large effect on the $\tilde{\Lambda}$ posterior, and reflect on whether they are robust to other modeling systematics. Finally, using the likelihood provided by this analysis, we perform full-end-to-end EOS inference with the GW and PSR likelihoods together, for the different BNS population prior models discussed above. As anticipated by our simple order-of-magnitude argument using only the posterior for $\tilde{\Lambda}$, we find the BNS population priors have a substantial effect on conclusions about the nuclear EOS.
III.2.1 Reanalysis of GW170817 alone
Figure 7 shows our results for the joint posterior in mass, aligned effective inspiral spin $\chi_{\rm eff}$, and tidal deformability $\tilde{\Lambda}$, showing both the posterior (contours) and marginal likelihood (points), obtained using different prior assumptions about the BNS population. All marginal likelihoods accumulated during multiple independent analyses of GW170817, performed with different intrinsic priors but with the same data handling and priors over extrinsic parameters are shown. Some of these prior assumptions by construction limit the range of spins and $q$ allowed, so sharp edges in posteriors associated with the prior constraints are expected and natural. For the fiducial analysis (black), the marginal distributions of mass, mass ratio, spin, and $\tilde{\Lambda}$ are consistent with previously published interpretations using comparable models, settings, and data [1, 2]. For comparison and to highlight the impact of astrophysically plausible prior assumptions on source parameter inferences, this figure also shows analyses in which we require either $q>0.9$, $\chi_{i,z}>0$, or both. As the likelihood contains strong correlated features, these constraints impact the posterior distribution of tidal deformability $\tilde{\Lambda}$. As seen in the $\tilde{\Lambda}-\chi_{eff}$ 2-d posterior, when applying only the $q>0.9$ condition the strong correlation between $\tilde{\Lambda}$ and $\chi_{eff}$ requires negative $\chi_{eff}$ for small $\tilde{\Lambda}$. As a result, by further requiring that $\chi_{i}>0$, we eliminate the possibility that $\tilde{\Lambda}$ can be small (blue). I.e., when applied individually, these two constraints have modest impact. However, when imposed simultaneously, even these modest constraints produce qualitative changes in the inferred tidal deformability.
Figure 7: GW inference Benchmark inference of properties of GW170817 using RIFT and IMRPhenomPv2_NRTidalv2. Color scale shows the inferred marginal likelihood, increasing from blue to red over a range of 15 in log-likelihood (with gray indicating values more than $15$ below the peak); 2d contours show the 90% credible intervals; and one-dimensional distributions show the one-dimensional marginal posterior distributions. Parameters shown are the detector-frame chirp mass (denoted here by ${\cal M}_{c}$), mass ratio $q=m_{2}/m_{1}$, aligned effective inspiral spin $\chi_{\rm eff}$, and tidal deformability $\tilde{\Lambda}$. Results for our fiducial analysis (in black) shown here are very similar to published results available through GWOSC [1, 2]. Colors indicate results derived using fiducial priors (black); $q>0.8$ (red); $\chi_{i,z}>0$ (green); and both (blue). The magenta curve shows results derived with a spin prior extending up to $|\chi_{i}|=0.4$.
Figure 8: GW inference assuming high mass ratio A reanalysis of GW170817 using RIFT and IMRPhenomPv2_NRTidalv2 while requiring $m_{2}/m_{1}<0.6$ As in Figure 7, the color scale shows the marginal likelihood and lines show one or two-dimensional marginal distributions. The solid black line adopts the fiducial spin prior $|\chi_{i}|<0.05$; the red line adopts a prior of similar form but allows much larger component spins $|\chi_{i}|<0.4$. Unlike the results presented in Figure 7, these investigations use a uniform prior on $\chi_{i,z}$, rather than the “z prior”, to ensure comprehensive coverage. Due to strong correlations between spin and mass ratio, the fiducial spin prior’s arbitrary upper limit strongly constrains the extent of the mass ratio distribution.
Figures 7 and 8 also illustrate inferences derived using even more extreme prior assumptions about the pertinent neutron star spin and mass ratio. For example, for the magenta curve provided in Figure 7, the analysis merely allowed for larger possible values of the NS maximum spin. Relaxing the upper limit on $\chi_{i}$ allows a lower mass for $m_{2}$ to remain consistent with the data [2], while increasing $\chi_{i}$ only incrementally. Though this single changed assumption allows for more extreme masses, in isolation this one relaxed assumption doesn’t change our qualitative interpretation; notably, the distribution of $\tilde{\Lambda}$ is nearly unchanged. By contrast, Figure 8 shows inferences derived assuming $m_{2}/m_{1}<0.6$, both adopting fiducial spin magnitudes (black) and larger NS spins (red). Notably, the likelihood includes comparable values for a wide range of mass ratios, including masses $m_{2}<1~{}{\rm M_{\odot}}$. Contrasting the likelihood (points) to the two posterior marginals (black and red) shows how strongly the arbitrary upper limit in our fiducial spin prior limits the mass ratio distribution. Both extreme-$q$ analyses nonetheless produce comparable $\tilde{\Lambda}$ posterior distributions. Due to correlations between mass ratio and $\tilde{\Lambda}$, the constraint $q<0.6$ causes the posterior distribution of $\tilde{\Lambda}$ to systematically decrease. In Appendix C, we show that the median $\tilde{\Lambda}$ steadily decreases versus an imposed maximum mass ratio $q$.
Figure 9: GW-derived NS radius Nominal distribution of $R_{1.4}$ derived from Eq. (10) and three GW-only inferences of GW170817 presented in Figures 7, 8: our fiducial analysis (blue); the comparable mass positive spin analysis (orange); and the high mass ratio analysis (green). Cumulative distribution is shown in solid lines and probability distribution is shown in lighter shade.
These GW-only reanalyses suggest that hidden astrophysical priors change $\tilde{\Lambda}$ by roughly a factor of 2, increasing the median $\tilde{\Lambda}$ (if we require comparable mass and positive aligned spins) or decreasing it (if we require $q<0.6$) respectively. However, because of the exponent relating $R_{1.4}$ and $\tilde{\Lambda}$, see Eq. (10), these factors of 2 propagate to a roughly 10% change in the inferred $R_{1.4}$, comparable to the statistical measurement errors usually presented for many of the individual measurements described earlier in Section II.3. Figure 9 shows the cumulative distribution of $R_{1.4}$ derived from this expression and three GW-only analyses: our fiducial analysis (blue); the comparable mass positive spin analysis (orange); and the high mass ratio analysis (green). Note, the bimodality in the probability distributions of the fiducial and the comparable-mass positive spin analyses arise due to bimodality in $\tilde{\Lambda}$ seen in Fig. 7.
We emphasize this simple discussion presents radius constraints directly in terms of conventional nonphysical GW priors on $\Lambda_{i}$, treating each as independent and uniformly distributed. These statistical errors are not informed by any EoS model families, nor even by the requirement that both NS have the same EoS. In addition, the jointly-uniform prior also produces a marginal prior distribution on $\tilde{\Lambda}$ that goes to zero as $\tilde{\Lambda}$ (i.e., $p(\tilde{\Lambda})\propto\tilde{\Lambda}$ at small $\tilde{\Lambda}$). To facilitate visualization and downstream calculation using posterior samples, other studies have reweighted the GW-only posterior distribution for example to be uniform in $\tilde{\Lambda}$ or $R_{1.4}$; see, e.g., Figure 11 of Ref. [2]. These two priors differ substantially: using Eq. (10), a uniform prior in $\tilde{\Lambda}$ (i.e., $P(<\Lambda)=\Lambda/\Lambda_{max}$) corresponds to a cumulative prior distribution that strongly favors the largest radii (i.e., $P(<R_{1.4})\propto R_{1.4}^{6}$). As our calculations employ a directly-interpolated marginal likelihood $L(X)$, not a smoothed posterior, we do not produce such an intermediate data product. Finally, even though the various model-agnostic GW priors used to construct the posterior distributions shown in Figs. 7, 8 and 9 strongly disfavor extremely small NS radii, the GW likelihood itself does not: very small $\Lambda_{i}$ are nearly as consistent with GW170817 observations as $\Lambda_{i}\simeq O(100)$.
III.2.2 Joint EOS inference using GW170817
@@
Gravitational wave measurements, currently favoring smaller radii, are by our prior’s construction more informative than NICER radius constraints. As noted in Figure 1 and the associated discussion, our EoS family and prior enforce a reasonably high radius for low-mass NS leading, through regularity of the mass-radius curve and a need for large maximum mass, to EoS models that favor comparable or larger radii for typical NS. For context, comparing Figures 2 and 1, our prior favors NS radii largely consistent with the radius of J0030, including a wide range between $9~{}{\rm km}$ and $16~{}{\rm km}$ for a fiducial $1.4~{}{\rm M_{\odot}}$ NS. By contrast, as surveyed in Figure 9, GW observations of GW170817 support typically smaller radii for a similar fiducial NS, with little support above $13~{}{\rm km}$ in a direct interpretation of our fiducial GW analyses. As a result and as seen in previous studies, a joint analysis incorporating both galactic pulsars and GW observations finds fiducial pulsar radii comparable to the intersection of these two limits. Figure 10 shows our fiducial calculation incorporating galactic pulsar measurements, GW170817 (with our fiducial prior), and the symmetry energy, following [45]. Using similar parameterizations and observational constraints as a previous study, we find comparable inferences for the nuclear equation of state [45]. Some small differences between our inferences and prior work are expected for technical reasons; for example, we present 90% quantiles in radius at each fixed NS mass, while previous and other works adopt a NS central density prior distribution; see the appendix for further discussion.
@@
That said, our GW analyses also demonstrate that hidden astrophysical priors have a substantial impact on the inferred NS radius. To characterize this effect, Figure 11 shows a one-dimensional scatterplot of $\ln{\cal E}$ versus $R_{1.4}$ using the three different GW analyses presented in Figure 9. These three analyses draw on identical RIFT marginal likelihood data, differing only in the assumed binary NS population model for mass ratio and spin. We see that the GW signal places an upper bound on the NS radius, and no strong lower bound (base). High mass ratio case (highSpin_lowQ) more strongly and more frequently penalizes higher NS radii, though some configurations with large radii are still highly likely; the overall likelihood shifts to the lowest radii. The comparable masses and positive $\chi_{i,z}$ (pos_highQ) strongly disfavours small radii and is bound on the higher radii showing most of the likelihood in a narrow range of radii on the higher end. This affirms (directly within the context of EoS inference) that astrophysical priors embedded in the GW-only analysis propagate directly into different conclusions about the nuclear EoS.
@@
The pulsar-only likelihoods represented in Figures 2 and 4 and the (multiple) GW-only likelihoods characterized by Figure 11 show that the pulsar and GW measurements may or may not agree, with the amount of tension depending strongly on hidden astrophysical priors. In one scenario, the pulsar/GW results may be in modest tension, consistent with previously presented work. In another scenario, the two may be in good agreement, favoring radii of $13~{}{\rm km}$, if we assume GW170817 is similar to galactic pulsars. Finally, the two approaches could have strong tension, with HESS/GW favoring small radii and NICER favoring large radii, if we assume the kilonova associated with GW170817 favors asymmetric NS binaries. As an example, Figure 12 shows an analysis comparable to Figure 10, excepting only the mass ratio prior ($q<0.6$) applied for GW170817. Despite adopting the same raw GW observation and inference strategy, the same galactic pulsar measurements, and the same symmetry energy, in this scenario we would infer a notably smaller radius for most pulsars, compared to previous work.
Figure 11: GW constraints for different astrophysical priors Scatterplot of $\ln{\cal E}$ versus $R_{1.4}$ evaluated for the same three astrophysical priors for NS binaries illustrated in Figure 9. As in that figure, this representation uses evaluations drawn from the $\Gamma$-spectral EoS. Top panel: Likelihood for the fiducial spin and mass ratio priors (base). Center panel: Likelihood for the high-mass ratio prior (highSpin_lowQ). Bottom panel: Likelihood for the comparable-mass and positive-$\chi_{i,z}$ prior (pos_highQ).
@@
Figure 12: Joint EoS constraint from GW, galactic pulsars, and symmetry energy, derived using $\Gamma$-spectral EoS parameterization, using three different prior assumptions about GW170817. The blue contours show results derived using the conventional 170817 prior (uniform masses); the orange requires $\chi_{i,z}>0$; and the green requires $q<0.6$. The analyses shown in blue have also been presented in Fig. 10. Top-left panel: 90% percentile pointwise confidence intervals for $p(\rho)$. Top-right panel: 90% percentile pointwise confidence interval for radius at each mass. Bottom panel: $\Gamma$-spectral parameter posterior distribution; we show marginal likelihood evaluations accumulated for the fiducial analysis (corresponding to the blue contours).